# Continuous-time Risk-sensitive Reinforcement Learning via Quadratic Variation Penalty

### v1

@@ 2.2 Reinforcement learning formulation
I now present the exploratory formulation in the sense of Wang et al., (2020) of the problem to be studied in this paper, which is regarded as the mathematical model for analyzing continuous-time RL. Mathematically, it means to generalize the classical problem to allow relaxed control (Fleming and Nisio, , 1984; Zhou, , 1992), where the actions are randomized generated from a probability distribution. Intuitively, what an agent can do is to try a (randomly generated) sequence of actions ${\boldsymbol{a}}=\{a_{s},t\leq s\leq T\}$ [^4] Arguably, it is not possible to sample continuously in practice, I keep to this continuous-time sequence for the ease of the development of algorithms. In principle, one can use simple process to approximate it (Szpruch et al., , 2024), or argue the existence via suitable measure-theory-based techniques (Sun, , 2006)., observe the resulting state process $X^{\boldsymbol{a}}=\{X_{s}^{\boldsymbol{a}},t\leq s\leq T\}$, and the realized reward $\{r(s,X_{s}^{\boldsymbol{a}},a_{s}),t\leq s\leq T\}$. Meanwhile, the agent gradually adjusts actions based on these observations. [^5] This procedure applies to both the offline and online settings. In the former, the agent can repeatedly try different sequences of actions over the same time period $[0,T]$ and record the corresponding state processes and payoffs. In the latter, the agent updates the actions as she goes, based on all the up-to-date historical observations. An agent can only do trial-and-error in the RL context because the agent has no knowledge about the environment (i.e. the functions $b,\sigma,r,h$) and cannot even form an HJB equation or any mathematical tool in the classical stochastic control methodology.
@@
When the expectation is taken with respect to only $X^{\boldsymbol{\pi}}$ under $\mathbb{P}$, it can be equivalently written as the expectation respect to $\tilde{X}^{\boldsymbol{\pi}}$ under $\mathbb{P}^{W}$. When the expectation involves $a^{\boldsymbol{\pi}}$ under $\mathbb{P}$, then it can be integrated out by conditioning on $X^{\boldsymbol{\pi}}$ using the policy $\boldsymbol{\pi}$ [^6] In Wang et al., (2020), this result is derived using the law-of-large-number type of argument to integrate over the randomization of $a$. More discussions on the relation between (5) and (6) can be found in Jia and Zhou, 2022b ; Jia and Zhou, (2023).
Based on the same spirit as in Wang et al., (2020), I add an entropy regularizer to the reward function to encourage exploration (represented by the stochastic policy), hence the risk-sensitive objective (2) becomes
@@
is an ($(\{\mathcal{F}_{s}^{X}\}_{s\geq t},\mathbb{P})$- (local) martingale; where $\langle J^{*^{\boldsymbol{\pi}}}\rangle$ is the QV of the process $J^{*}(s,X_{s}^{\boldsymbol{\pi}};\epsilon)$. Then $\boldsymbol{\pi}^{*}$ is the optimal solution to (7) and $J^{*}(t,x;\epsilon)$ is the optimal value function.
@@ 3.1 Definition of q-function
More precisely, for a given policy $\boldsymbol{\pi}$, and $(t,x,a)\in[0,T)\times\mathbb{R}^{d}\times\mathcal{A}$, consider a “perturbed” policy of $\boldsymbol{\pi}$ as follows: It takes the action $a\in\mathcal{A}$ on $[t,t+\Delta t)$ where $\Delta t>0$, and then follows $\boldsymbol{\pi}$ on $[t+\Delta t,T]$. Then the advantage function is
@@
×$\times\left[r(s,X_{s}^{a},a)+\mathcal{L}^{a}J(s,X_{s}^{a};\boldsymbol{\pi};\epsilon)+\frac{\epsilon}{2}|\sigma(s,X_{s}^{a},a)^{\top}\frac{\partial J}{\partial x}(s,X_{s}^{a};\boldsymbol{\pi};\epsilon)|^{2}\right]\mathrm{d}s\Big{|}X_{t}^{a}=x\Bigg{]}$
@@ 3.2 Martingale characterization of the optimal q-function
    is an ($(\{\mathcal{F}_{s}\}_{s\geq t},\mathbb{P})$-martingale, where $\{{X}_{s}^{\boldsymbol{\pi}},t\leq s\leq T\}$ is the solution to (5) under $\boldsymbol{\pi}$ with ${X}_{t}^{\boldsymbol{\pi}}=x$.
@@
    Suppose stronger regularity condition $\mathbb{E}\left[e^{\frac{\epsilon^{2}}{2}\langle\widehat{J^{*}}^{\boldsymbol{\pi}}\rangle(T)}\right]<\infty$ for any $\boldsymbol{\pi}\in\boldsymbol{\Pi}$ holds. If there exists one $\boldsymbol{\pi}\in\boldsymbol{\Pi}$ such that for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, (15) is an ($(\{\mathcal{F}_{s}\}_{s\geq t},\mathbb{P})$-martingale, then $\widehat{J^{*}}$ and $\widehat{q^{*}}$ are respectively the optimal value function and the optimal q-function. Moreover, in this case, $\widehat{\boldsymbol{\pi}^{*}}(a|t,x)=\exp\{\frac{1}{\lambda}\widehat{q^{*}}(t,x,a)\}$ is the optimal policy.
@@ 3.3 Risk-sensitive q-Learning
Note that based on the martingale characterization of the optimal q-function, it is possible to device various algorithms that either aim to minimize a suitable loss function, or to solve a system of moment conditions, as evidenced in Jia and Zhou, (2023). Then one may use either stochastic gradient decent or stochastic approximation algorithm to numerically find the solutions.
@@
    end while $\triangleright$ Off-policy case Obtain one observation $\{a_{t_{k}},r_{t_{k}},x_{t_{k+1}}\}_{k=0,\cdots,K-1}\cup\{x_{t_{K}},h(x_{t_{K}})\}=\textit{Observation}(\Delta t)$. $\triangleright$ After getting a trajectory For every $i=0,1,\cdots,K-1$, compute and store test functions $\xi_{t_{i}}=\boldsymbol{\xi}(t_{i},x_{t_{0}},\cdots,x_{t_{i}},a_{t_{0}},\cdots,a_{t_{i}})$, $\zeta_{t_{i}}=\boldsymbol{\zeta}(t_{i},x_{t_{0}},\cdots,x_{t_{i}},a_{t_{0}},\cdots,a_{t_{i}})$. Compute
@@ 4.1 Contrast with policy gradient
-- $\lambda\log\boldsymbol{\pi}^{\phi}(a_{s}^{\boldsymbol{\pi}^{\phi}}|s,X_{s}^{\boldsymbol{\pi}^{\phi}})\mathrm{d}s+\frac{\epsilon}{2}\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}}\rangle(s)\bigg{]}+\epsilon\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}},G^{\boldsymbol{\pi}^{\phi}}\rangle(s)\mid X_{t}^{\boldsymbol{\pi}^{\phi}}=x\Bigg{]}.$
In the above expression, note that $\mathrm{d}J^{\boldsymbol{\pi}^{\phi}}_{s}+r(s,X_{s}^{\boldsymbol{\pi}^{\phi}},a_{s}^{\boldsymbol{\pi}^{\phi}})\mathrm{d}s-\lambda\log\boldsymbol{\pi}^{\phi}(a_{s}^{\boldsymbol{\pi}^{\phi}}|s,X_{s}^{\boldsymbol{\pi}^{\phi}})\mathrm{d}s$ is the usual “temporal difference” (TD) term, and it is augmented by the QV $\frac{\epsilon}{2}\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}}\rangle(s)$ due to the QV penalty. It is remarkable that the policy gradient is no longer the gradient of the log-likelihood of the policy multiplied by the (augmented) TD error, instead, an extra term arises because of the nonlinear nature. Furthermore, this extra term $\epsilon\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}},G^{\boldsymbol{\pi}^{\phi}}\rangle(s)$ is the cross-variation between the value process and its gradient process, and remains uncomputable directly. Therefore, it is difficult to obtain an unbiased estimate of the policy gradient in the risk-sensitive RL. In contrast, q-learning does not have this problem since the universal martingale property. Example 1 provides a simple illustration of the bias caused by ignoring this extra term in estimating the policy gradient, and such bias is proportional to the risk sensitivity coefficient $\epsilon$.
@@ 4.2 Ergodic tasks
The q-function in such ergodic task can be similarly defined as in Definition 2 except that the associated value function $J(\cdot;\boldsymbol{\pi};\epsilon)$ becomes time-invariant, and is only unique up to a constant. In addition, there is a number $\beta(\boldsymbol{\pi};\epsilon)$ that is associated with the policy. This number stands for the objective function value (the long-term average) and it is also part of the solutions.
@@
$\lim_{T\to\infty}\mathbb{E}\left[e^{\frac{\epsilon^{2}}{2}\langle\widehat{J^{*}}^{\boldsymbol{\pi}}\rangle(T)}\right]<\infty,\text{ for any }\boldsymbol{\pi}\in\boldsymbol{\Pi},\ \int_{\mathcal{A}}\exp\{\frac{1}{\lambda}\widehat{q^{*}}(x,a)\}\mathrm{d}a=1,\;\;\forall x\in\mathbb{R}^{d}.$ (18)
Moreover, assume there exists $\delta>0$ such that $\lim_{T\to\infty}\frac{1}{T}\log\mathbb{E}^{\mathbb{P}}\left[e^{-\epsilon(1+\delta)\widehat{J^{*}}(X_{T}^{\boldsymbol{\pi}})}\right]=0$ under any probability measure $\hat{\mathbb{P}}$ that is equivalent to the original probability $\mathbb{P}$, for all $\boldsymbol{\pi}\in\boldsymbol{\Pi}$.
If there exists one $\boldsymbol{\pi}\in\boldsymbol{\Pi}$ such that for all initial state $x\in\mathbb{R}^{d}$, the following process
@@
is an ($(\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$-martingale. Then $\widehat{J^{*}}$, $\widehat{q^{*}}$, $\widehat{\beta^{*}}$ are respectively the optimal value function, the optimal q-function, and the optimal value. Moreover, in this case, $\widehat{\boldsymbol{\pi}^{*}}(a|x):=\exp\{\frac{1}{\lambda}\widehat{q^{*}}(x,a)\}$ is the optimal policy.
@@ 5.1 Merton’s investment problem with power utility
The problem for an RL agent is, to find the optimal policy to this problem without the knowledge of $\mu,\sigma$. ($\gamma$ is assumed to be known by the RL agent) At the first glance, the problem belongs to standard, non-risk-sensitive RL, and prior studies have been done to attack this problem directly. However, they have been largely restricted to log-utility, e.g., Dai et al., 2023a ; Jiang et al., (2022) due to the tractability of the associated exploratory stochastic control problems, which correspond to the limit of $\gamma\to 1$. Even for general power utilities, the associated exploratory stochastic control problems have to be specifically design to guarantee well-posedness and tractability, see discussions in Dai et al., 2023b .
However, this problem can be naturally embedded into a risk-sensitive RL framework by utilizing the transformation: $U(x)=\frac{1}{1-\gamma}e^{(1-\gamma)\log x}$, so that a power utility can be reviewed as the risk-sensitive counterpart of the log-returns of the portfolio, where $10\neq 1-\gamma<$ 1 is the risk sensitivity coefficient. This connection has been noticed in the mean-variance analysis for the log-returns in Dai et al., (2021) due to the close relation between a risk-sensitive problem and the mean-variance problem and their applications in portfolio management (Bielecki and Pliska, , 1999).

### v2

@@ 2.2 Reinforcement learning formulation
I now present the exploratory formulation in the sense of Wang et al., (2020) of the problem to be studied in this paper, which is regarded as the mathematical model for analyzing continuous-time RL. Mathematically, it means to generalize the classical problem to allow relaxed control, where the actions are randomized generated from a probability distribution. Intuitively, what an agent can do is to try a (randomly generated) sequence of actions ${\boldsymbol{a}}=\{a_{s},t\leq s\leq T\}$ [^4] Arguably, it is not possible to sample continuously in practice, I keep to this continuous-time sequence for the ease of the development of algorithms. In principle, one can use simple process to approximate it (Szpruch et al., , 2024)., observe the resulting state process $X^{\boldsymbol{a}}=\{X_{s}^{\boldsymbol{a}},t\leq s\leq T\}$, and the realized reward $\{r(s,X_{s}^{\boldsymbol{a}},a_{s}),t\leq s\leq T\}$. Meanwhile, the agent gradually adjusts actions based on these observations. [^5] This procedure applies to both the offline and online settings. In the former, the agent can repeatedly try different sequences of actions over the same time period $[0,T]$ and record the corresponding state processes and payoffs. In the latter, the agent updates the actions as she goes, based on all the up-to-date historical observations. An agent can only do trial-and-error in the RL context because the agent has no knowledge about the environment (i.e. the functions $b,\sigma,r,h$) and cannot even form an HJB equation or any mathematical tool in the classical stochastic control methodology.
@@
When the expectation is taken with respect to only $X^{\boldsymbol{\pi}}$ under $\mathbb{P}$, it can be equivalently written as the expectation respect to $\tilde{X}^{\boldsymbol{\pi}}$ under $\mathbb{P}^{W}$. When the expectation involves $a^{\boldsymbol{\pi}}$ under $\mathbb{P}$, then it can be integrated out by conditioning on $X^{\boldsymbol{\pi}}$ using the policy $\boldsymbol{\pi}$. [^6] In Wang et al., (2020), this result is derived using the law-of-large-number type of argument to integrate over the randomization of $a$. More discussions on the relation between (5) and (6) can be found in Jia and Zhou, 2022b ; Jia and Zhou, (2023).
Based on the same spirit as in Wang et al., (2020), I add an entropy regularizer to the reward function to encourage exploration (represented by the stochastic policy), hence the risk-sensitive objective (2) becomes [^7] When $\lambda=0$, (7) reduces to the classical risk-sensitive control problem (2). For the general non-risk-sensitive problems, the entropy-regularized problem approximates the classical problem by the order of $O(-\lambda\log\lambda)$, see Tang et al., (2022). However, whether the same order applies to the risk-sensitive problem still remains an open question. For the special case of Merton’s investment problem studied in Section 5, the connection between the entropy-regularized problem and the classical problem becomes clear because the closed-form solution to the entropy-regularized problem can be found in Appendix A.1. In this case, the gap is indeed $\frac{\lambda}{2}\log\frac{2\pi\lambda}{\gamma\sigma^{2}}$.
@@
is an $(\{\mathcal{F}_{s}^{X}\}_{s\geq t},\mathbb{P})$- (local) martingale; where $\langle J^{*^{\boldsymbol{\pi}}}\rangle$ is the QV of the process $J^{*}(s,X_{s}^{\boldsymbol{\pi}};\epsilon)$. Then $\boldsymbol{\pi}^{*}$ is the optimal solution to (7) and $J^{*}(t,x;\epsilon)$ is the optimal value function.
@@ 3.1 Definition of q-function
More precisely, for a given policy $\boldsymbol{\pi}$, and $(t,x,a)\in[0,T)\times\mathbb{R}^{d}\times\mathcal{A}$, consider a “perturbed” policy of $\boldsymbol{\pi}$ as follows: It takes the action $a\in\mathcal{A}$ on $t,t+\Delta t)$ where $\Delta t>0$, and then follows $\boldsymbol{\pi}$ on $t+\Delta t,T]$. Then the advantage function is
@@
$\times\left[r(s,X_{s}^{a},a)+\mathcal{L}^{a}J(s,X_{s}^{a};\boldsymbol{\pi};\epsilon)+\frac{\epsilon}{2}|\sigma(s,X_{s}^{a},a)^{\top}\frac{\partial J}{\partial x}(s,X_{s}^{a};\boldsymbol{\pi};\epsilon)|^{2}\right]\mathrm{d}s\Big|X_{t}^{a}=x\Bigg]$
@@ 3.2 Martingale characterization of the optimal q-function
    is an $(\{\mathcal{F}_{s}\}_{s\geq t},\mathbb{P})$-martingale, where $\{{X}_{s}^{\boldsymbol{\pi}},t\leq s\leq T\}$ is the solution to (5) under $\boldsymbol{\pi}$ with ${X}_{t}^{\boldsymbol{\pi}}=x$.
@@
    Suppose stronger regularity condition $\mathbb{E}\left[e^{\frac{\epsilon^{2}}{2}\langle\widehat{J^{*}}^{\boldsymbol{\pi}}\rangle(T)}\right]<\infty$ for any $\boldsymbol{\pi}\in\boldsymbol{\Pi}$ holds. If there exists one $\boldsymbol{\pi}\in\boldsymbol{\Pi}$ such that for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, (15) is an $(\{\mathcal{F}_{s}\}_{s\geq t},\mathbb{P})$-martingale, then $\widehat{J^{*}}$ and $\widehat{q^{*}}$ are respectively the optimal value function and the optimal q-function. Moreover, in this case, $\widehat{\boldsymbol{\pi}^{*}}(a|t,x)=\exp\{\frac{1}{\lambda}\widehat{q^{*}}(t,x,a)\}$ is the optimal policy.
@@
Note that, the conditions required in ($ii$) are imposed jointly on the policy $\boldsymbol{\pi}$, the candidate value function and q-function $\widehat{J^{*}},\widehat{q^{*}}$; by ($i$), if the candidate value function and q-function are the optimal value function and q-function, then this policy $\boldsymbol{\pi}$ can be any arbitrary policy. Therefore, as long as the original problem (7) admits an optimal policy, then we do not need to impose additional restrictions.
@@ 3.3 Risk-sensitive q-Learning
Note that based on the martingale characterization of the optimal q-function, it is possible to devise various algorithms that either aim to minimize a suitable loss function, or to solve a system of moment conditions, as evidenced in Jia and Zhou, (2023). Then one may use either stochastic gradient decent or stochastic approximation algorithm to numerically find the solutions.
@@
[…]
@@ 4.1 Contrast with policy gradient
- $\lambda\log\boldsymbol{\pi}^{\phi}(a_{s}^{\boldsymbol{\pi}^{\phi}}|s,X_{s}^{\boldsymbol{\pi}^{\phi}})\mathrm{d}s+\frac{\epsilon}{2}\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}}\rangle(s)\bigg]$
$+$ $\epsilon\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}},G^{\boldsymbol{\pi}^{\phi}}\rangle(s)\mid X_{t}^{\boldsymbol{\pi}^{\phi}}=x\Bigg].$
In the above expression, note that $\mathrm{d}J^{\boldsymbol{\pi}^{\phi}}_{s}+r(s,X_{s}^{\boldsymbol{\pi}^{\phi}},a_{s}^{\boldsymbol{\pi}^{\phi}})\mathrm{d}s-\lambda\log\boldsymbol{\pi}^{\phi}(a_{s}^{\boldsymbol{\pi}^{\phi}}|s,X_{s}^{\boldsymbol{\pi}^{\phi}})\mathrm{d}s$ is the usual “temporal difference” (TD) term, and it is augmented by the QV $\frac{\epsilon}{2}\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}}\rangle(s)$ due to the QV penalty. It is remarkable that the policy gradient is no longer the gradient of the log-likelihood of the policy multiplied by the (augmented) TD error, instead, an extra term arises because of the nonlinear nature. Furthermore, this extra term $\epsilon\mathrm{d}\langle J^{\boldsymbol{\pi}^{\phi}},G^{\boldsymbol{\pi}^{\phi}}\rangle(s)$ is the cross-variation between the value process and its gradient process, and remains uncomputable directly. Therefore, it is difficult to obtain an unbiased estimate of the policy gradient in the risk-sensitive RL. In contrast, q-learning does not have this problem due to the universal martingale property. Example 1 provides a simple illustration of the bias caused by ignoring this extra term in estimating the policy gradient, and such bias is proportional to the risk sensitivity coefficient $\epsilon$.
@@ 4.2 Ergodic tasks
The q-function in such ergodic task can be similarly defined as in Definition 2 except that the associated value function $J(\cdot;\boldsymbol{\pi};\epsilon)$ becomes time-invariant, and is only unique up to a constant. In addition, there is a number $\beta(\boldsymbol{\pi};\epsilon)$ that is associated with the policy. This number stands for the objective function value (the long-term average) and it is also part of the solutions. Moreover, since the ergodic tasks focus on the long-term average performance, we restrict the SDE (6) to be time-homogeneous and also restrict the admissible policies to time-invariant ones.
Definition 3.
A time-invariant policy $\boldsymbol{\pi}=\boldsymbol{\pi}(\cdot|\cdot)$ is called admissible if
  1. (i)
    $\boldsymbol{\pi}(\cdot|x)\in\mathcal{P}(\mathcal{A})$, $\operatorname{supp}\boldsymbol{\pi}(\cdot|x)=\mathcal{A}$ for every $x\in\mathbb{R}^{d}$, and $\boldsymbol{\pi}(a|x):(x,a)\in\mathbb{R}^{d}\times\mathcal{A}\mapsto\mathbb{R}$ is measurable;
  2. (ii)
    the SDE (6) admits a unique weak solution (in the sense of distribution) for any initial $(t,x)\in\mathbb{R}_{+}\times\mathbb{R}^{d}$;
  3. (iii)
    $\int_{\mathcal{A}}|r(x,a)-\gamma\log\boldsymbol{\pi}(a|x)|\boldsymbol{\pi}(a|x)\mathrm{d}a\leq C(1+|x|^{\mu})$, $\forall(t,x)$ where $C>0$ and $\mu\geq 1$ are constants;
  4. (iv)
    $\boldsymbol{\pi}(a|x)$ is continuous in $x$ and uniformly Lipschitz continuous in $x$ in the total variation distance, i.e., for each fixed $a$, $\int_{\mathcal{A}}|\boldsymbol{\pi}(a|x)-\boldsymbol{\pi}(a|x^{\prime})|\mathrm{d}a\to 0$ as $x^{\prime}\to x$, and there is a constant $C>0$ independent of $a$ such that
    $\int_{\mathcal{A}}|\boldsymbol{\pi}(a|x)-\boldsymbol{\pi}(a|x^{\prime})|\mathrm{d}a\leq C|x-x^{\prime}|,\;\;\forall x,x^{\prime}\in\mathbb{R}^{d}.$
The collection of admissible time-invariant policies is denoted by $\boldsymbol{\Pi}_{s}$.
@@
$\lim_{T\to\infty}\mathbb{E}\left[e^{\frac{\epsilon^{2}}{2}\langle\widehat{J^{*}}^{\boldsymbol{\pi}}\rangle(T)}\right]<\infty,\text{ for any }\boldsymbol{\pi}\in\boldsymbol{\Pi}_{s},\ \int_{\mathcal{A}}\exp\{\frac{1}{\lambda}\widehat{q^{*}}(x,a)\}\mathrm{d}a=1,\;\;\forall x\in\mathbb{R}^{d}.$ (18)
Moreover, assume there exists $\delta>0$ such that $\lim_{T\to\infty}\frac{1}{T}\log\mathbb{E}^{\hat{\mathbb{P}}}\left[e^{-\epsilon(1+\delta)\widehat{J^{*}}(X_{T}^{\boldsymbol{\pi}})}\right]=0$ under any probability measure $\hat{\mathbb{P}}$ that is equivalent to the original probability $\mathbb{P}$, for all $\boldsymbol{\pi}\in\boldsymbol{\Pi}_{s}$.
If there exists one $\boldsymbol{\pi}\in\boldsymbol{\Pi}_{s}$ such that for all initial state $x\in\mathbb{R}^{d}$, the following process
@@
is an $(\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$-martingale. Then $\widehat{J^{*}}$, $\widehat{q^{*}}$, $\widehat{\beta^{*}}$ are respectively the optimal value function, the optimal q-function, and the optimal value. Moreover, in this case, $\widehat{\boldsymbol{\pi}^{*}}(a|x):=\exp\{\frac{1}{\lambda}\widehat{q^{*}}(x,a)\}$ is the optimal policy.
@@ 5.1 Merton’s investment problem with power utility
The problem for an RL agent is, to find the optimal policy to this problem without the knowledge of $\mu,\sigma$. ($\gamma$ is assumed to be known by the RL agent) At the first glance, the problem belongs to standard, non-risk-sensitive RL, and prior studies have studied this problem directly. However, they have been largely restricted to log-utility, e.g., Dai et al., (2023); Jiang et al., (2022) due to the tractability of the associated exploratory stochastic control problems, which correspond to the limit of $\gamma\to 1$. Even for general power utilities, the associated exploratory stochastic control problems have to be specifically design to guarantee well-posedness and tractability, see discussions in Dai et al., (2026).
However, this problem can be naturally embedded into a risk-sensitive RL framework by utilizing the transformation: $U(x)=\frac{1}{1-\gamma}e^{(1-\gamma)\log x}$, so that a power utility can be reviewed as the risk-sensitive counterpart of the log-returns of the portfolio, where $0\neq 1-\gamma<1$ is the risk sensitivity coefficient. This connection has been noticed in the mean-variance analysis for the log-returns in Dai et al., (2021) due to the close relation between a risk-sensitive problem and the mean-variance problem and their applications in portfolio management (Bielecki and Pliska, , 1999).